# Thinking twice inside the box: is Wigner’s friend really quantum?

### v1

@@ III.1 Restrictions on descriptions of physical systems and agents
Proponents of the Lewisian ‘worm view’, on the other hand, might say that the type of uncertainty that, for example, Freya faces at the beginning of Thought Experiment 2 is nothing but instantaneous self-locating uncertainty: Freya does not know which spacetime worm she is (the one passing through the green room, or the other one passing through the blue room at future spacetime points). At least in the context of classical physics, Lewisians might argue that their view suggests a natural ‘division of labour’: epistemology or the philosophy of mind would be the fields to consult if Freya would like to obtain guidance in the face of her self-locating uncertainty, whereas physics can inform her (and us) about which types of spacetime worms exist, or which ones are probable, according to the laws of (statistical) mechanics. Restriction A would then be an unavoidable consequence for all theories that we deem physical. However, this conclusion becomes problematic in light of quantum theory: according to some Everettian interpretations, quantum probabilities are of a similar kind to the self-locating uncertainty probabilities (according to Lewisians) in branching scenarios like Thought Experiment 2, and yet, quantum theory as a physical theory has certainly something to say about those probabilities.
@@
In Subsection III.3, we will argue that Restriction A can be considered an essential ingredient of another puzzle in the foundations of physics and philosophy: the Boltzmann brain problem. Before we turn to this, let us discuss in more detail how Restrictions P and A are related in quantum theory.
III.2 Quantum physics: Restriction A from Restriction P
In our terminology, EWF experiments are implemented to lift Bell violations (demonstrations of Restriction P under the assumptions of Local Causality and No Superdeterminism) to LF violations (demonstrations of Restriction A under the assumptions of Locality and No Superdeterminism). This can be seen in Bong et al. [7], who show that a violation of an LF inequality always implies the violation of Bell inequality. In this sense, we can interpret the results in [7, 24] as grounding Restriction A in Restriction P via the identification of agents with physical systems.
@@ III.2 Quantum physics: Restriction A from Restriction P
The converse is not true in general: violating an LF inequality is a mathematically strictly stronger fact (and hence a metaphysically strictly more dramatic fact) than the violation of a Bell inequality. However, it is instructive to consider the results of [67], who have shown that all Bell violations can be lifted to some kind of LF violation. In this sense, Restriction P, implemented via Bell nonlocality, implies Restriction A: once we have the former, we can construct an instance of the latter. This is shown by considering sequential EWFSs. A sequential EWFS is a scenario in which the superobserving parties (in this case, just Alice) may unitarily reverse the lab and ask their friend to measure in a different basis multiple times prior to ending the experiment. After each measurement performed by Charlie, Alice chooses one of two settings $x_{i}\in\{0,1\}$, where $1\leq i<R$. If $x_{i}=0$ and $i<R$, she reverses the lab and asks Charlie to perform a new measurement. Or if $x_{i}=1$, she asks Charlie his outcome. For the $R$th iteration, after $R-1$ potential reversals, she measures the particle directly for $x_{R}=0$, or asks Charlie’s outcome for $x_{R}=1$. It is shown that for any two-party scenario $\mathcal{S}=(\mathcal{A},\mathcal{B},\mathcal{X},\mathcal{Y})$, the violation of a Bell inequality within a EWFS implies the violation of a LF inequality, i.e.
@@
where $\mathbb{L}\mathbb{D}$ denotes the polytope describing behaviours that satisfy local determinism (Bell correlations) and $\mathbb{S}\mathbb{W}$ those that satisfy the LF assumptions for a sequential EWFS, with $R=|\mathcal{X}|-1$. We will look now at the precise definition of these sets.
@@
A proof of equation (2) is presented in [67]. The first, more established direction is that probabilities in $\mathbb{L}\mathbb{D}(\mathcal{S})$ must also be contained within $\mathbb{S}\mathbb{W}(\mathcal{S}_{R,0})$. This had already been observed in [7], since the LF set of correlations is a strict superset of the LHV set. In our terminology, this is to say that an instance of Restriction A (i.e. a LF violation) entails an instance of Restriction P (i.e. that it does not satisfy a local hidden variable model). An intuition for the opposite direction can be obtained by noticing that AOE” implies both AOE’ and Predetermination’. First, AOE” implies AOE’ for $p(\tilde{a}b|\tilde{x}y)=\sum_{\overleftrightarrow{c}}p(\tilde{a}b\overleftrightarrow{c}|\tilde{x}y)$, when $\tilde{a}$ and $\tilde{x}$ are relabelled $a$ and $x$. Second, AOE” implies Predetermination’, as one can define a hidden variable $\zeta$ containing $\overleftrightarrow{c}$ and $j$, where $j$ labels the extreme points of $\mathbb{N}\mathbb{S}((\mathcal{A},\mathcal{B},R+1,\mathcal{Y}))$. Charlie’s outcomes $\overleftrightarrow{c}$ function as deterministic hidden variables for all but one of the settings chosen by Alice, whilst the distribution over $j$ corresponds to the convex mixture of the different deterministic non-signalling behaviours that are realised in the experiment. Intuitively, it is essentially the claim of the absoluteness of Charlie’s outcomes $\overleftrightarrow{c}$ that lets them play the role of a set of hidden variables. For details, please see the proof of [67]. In the terminology we introduce here: if Alice’s subsystem is an instance of Restriction P (as witnessed by the violation of Bell-type inequalities when considered together with Bob’s subsystem), then it can also be used to observe an instance of Restriction A (i.e. a LF violation).
The equivalence expressed by equation (2) demonstrates that all Bell violations can be lifted to some kind of LF violation. In other words, all instances of Restriction P arising from Bell violations can be modified to also yield demonstrations of Restriction A for quantum theory via an EWF scenario, when one assumes Local Agency. It is also interesting to note that, whilst the quantum thought experiments above all involve instances of violations of Bell-type inequalities, one could also begin with contextuality as an example of Restriction P, and lift this to examples of Restriction A. This is done, for example, in Refs. [69, 70].
III.3 Restriction A beyond quantum foundations: should you believe you are a Boltzmann brain?
We contend that further puzzles in physics and philosophy concerning identity and first-person experience can be reduced to Restriction A. In this subsection, we discuss one such example for which we argue that this is the case: the Boltzmann brain (BB) problem. For a brief introduction to this problem, see Appendix V.5. Here, we will be even more brief, and discuss only a schematic version ignoring all details that are irrelevant for our discussion. Our exposition will mainly follow Carroll’s work [71].
@@ III.3 Restriction A beyond quantum foundations: should you believe you are a Boltzmann brain?
In this paper, we have carefully avoided discussing situations in which agents wonder who or where they are, and we have defined the notion of Restriction A in terms of an agent’s observations. To connect the BB discussion to the one in earlier sections, let us therefore reformulate this statement in a way that makes it more operational:
(***) If Freya lives in a BB-dominated universe, then she should expect that she will probably make a BB observation soon (such as seeing the microwave background being replaced by thermal background).
We argue that this reformulation captures more carefully the empirical claim that underlies the standard argument (S) above: after all, we use (what we think are) our observations to conclude that we are not BBs. Moreover, quantum physics motivates some caution to ground physical or metaphysical claims in concrete observations as far as possible. For example, the question of whether a photon in an interferometer is in the left or the right arm cannot be given any immediate meaning unless it is operationalised in terms of an empirical observation (say, by asking whether a detector in the left arm does or does not click). Therefore, statement (***) is a more careful formulation than the one preceding it.
Statement (***) can also be understood as grounding (C), cognitive instability, in a more operational way: if we have randomly fluctuated into existence, then our future observations will probably be uncorrelated with our beliefs and presumed knowledge. In other words: assuming the cosmological model of a BB-dominated universe, (***) implies that all our future observations would be uncorrelated with all we know, including this model and its predictions, which in turn should undermine our trust in the model.
However, statement (***) is far from obvious. It relies on something like the following unspoken assumption:
@@
However, the statement of whether Freya will soon make a BB observation (more succinctly and imprecisely, whether “she” is a BB) is not a statement about facts of the world. As such, we cannot use our physical theories directly to assign a probability $p$ to it. More concretely, physics itself, i.e. our universe’s laws and mechanisms, cannot determine its truth value, or the chances of it becoming true. Even if there are laws of the universe that can be phrased in statistical language in the regime of approximation that is relevant for all copies of Freya, these laws will in general fail to imply directly an assignment of probability $p$. The vast probability space describing all relevant facts of the world, i.e. whatever classical facts will present themselves to you and your fellow physicists according to your joint choice of Heisenberg cut, will contain $N$ random variables (one for each representation or realisation of Freya), $M$ of them for BBs and $N-M$ of them for non-BBs. But if You, the reader, are “Freya”, defined by your locally available information and nothing else, then you cannot identify any random variable on this space that would describe You. Hence, it is impossible to obtain a marginal distribution from a joint distribution of “facts of the world” that would correspond to the correct probability assignment of what You should believe to observe next. This is an instance of Restriction A.
We can see the claim above (that one should assign probability $M/N$) as a particularly intuitive attempt to respond to Restriction A – that is, to follow Elga’s Principle of Indifference [40] in assigning a uniform probability over all realisations. However, the situation described by Elga (an entertaining scenario involving “Dr. Evil” summarised in Appendix V.4) involves only very few (two) copies of an agent, with the property that the future observations of each copy do not dramatically differ from each other in their information-theoretic properties. It is unclear though whether (or, at the very least, not self-evident that) the uniform distribution is what we should assign in all other cases. In particular, the indifference principle must fail if there are infinitely many copies of Freya. Unrealistic as this may seem, we should not exclude this case a priori, since we would like to construct a principled way to respond to Restriction A that would also work in a literally infinite universe. Furthermore, Everettians in particular may want to admit that (branch or copy) counting is not always what one should do, but other structural features should impact probability assignments [72]. Before describing an example of what could potentially replace the indifference principle, let us acknowledge that the above statement (***) relies crucially on this assumption, and hence:
@@
We cannot argue that the prediction of a Boltzmann Brain-dominated universe invalidates a cosmological model, unless we respond to Restriction A in a particular way [^3] The notion of “agent” is here somewhat different from that of the previous thought experiments, but still in accordance with Definition 2. Namely, in our thought experiments and in WF-type experiments, we can at least in its initial stage identify the agent with a fixed, uniquely identified material object. Here, however, the agent is defined as a user of a physical theory that happens to be in a specific local configuration, without specifying which specific piece of matter corresponds to it, or whether (for example) it is itself a Boltzmann brain or not. Still, in line with Definition 2, every reader can in principle imagine putting themselves in the shoes of this element..
@@
This is the main point we would like to make in this section: Restriction A is relevant beyond the thought experiment of Wigner’s friend. Its reappearance in other scenarios with direct relevance to physics should be regarded as an argument against an attitude of methodological ignorance or of complete deference to other fields of inquiry such as the Philosophy of Mind.
@@
Now, as demonstrated in [63], if the above prescription is assumed to define the chances of Freya’s future observations, then this particular way to respond to Restriction A will lead to a consequence that is in some sense the opposite to the one drawn via indifference: it implies that Freya should not expect to make BB-observations soon, even if she lives in a BB-dominated universe. In a nutshell, the reason for this is that the BB-observation of thermal, uncorrelated radiation makes the resulting data algorithmically uncorrelated with Freya’s previous observations, which makes $M(y|x)$ small — crucially, this value becomes much larger if $y$ describes “business as usual on Earth”. We refer the interested reader to [63] for details. This demonstrates the sensitivity of the usual BB argumentation to the choice of “self-location” measure, i.e. to the way in which we respond to Restriction A.
@@ IV Conclusions
In this article, we have presented a property “Restriction A”, which we argue is a feature of classical, quantum and indeed more general physical theories: essentially, that a physical theory cannot always provide a probabilistic description of the observations of all agents. We argue that the violation of a Local Friendliness inequality in Extended Wigner’s friend (EWF) scenarios proves a particularly dramatic instance of Restriction A for quantum theory, and all empirically adequate future theories, if one assumes Locality and No Superdeterminism, but that this should be seen amongst a broader class of puzzles that demonstrate the feature. In particular, we have presented a number of thought experiments involving the duplication of agents, for which classical physics provides no means of assigning outcome probabilities for the users of the theory. Moreover, even supplementing classical physics with some probability rule that supplies these predictions will inevitably lead to inconsistent predictions amongst agents (i.e. to a theory that is still subject to Restriction A), unless this probability rule disregards elements of the setup that are at least intuitively essential. Furthermore, we have shown that Restriction A is also at the heart of cosmology’s Boltzmann brain (BB) problem. In particular, the question of whether cosmological models that predict a BB-dominated universe should be disregarded a priori cannot be answered without reacting to Restriction A in a particular way.
@@
We take it that the ultimate importance of EWF and similar thought experiments lie in revealing a crucial methodological restriction of our current physical theories: that they typically only provide (probabilistic) predictions for situations in which these predictions can be intersubjectively tested by external observers. However, there are situations such as EWF scenarios for which potential predictions could not be intersubjectively, but privately tested: we could certainly imagine putting ourselves into the shoes of the Friend and testing our predictions for what we will observe, but our theories do not always tell us what to expect in situations of this kind. More broadly, and more conceptually, we do not always know how to tackle the epistemological barrier in asking: what is it like to be a Friend? [74] How can I fully understand other perspectives, and are they even compatible with my own? In analogy with Nagel’s argument, our claim here is that external facts supplied by our physical theories cannot provide a complete picture of internal experiences and predictions. Crucially, this observation does not only refer to the obvious restriction of not knowing what it feels like to be a Friend, but even of asking whether the Friend will hear a detector click or not. It may be tempting to ignore such questions as “unphysical”, since their answers cannot be verified by experiments performed by external observers. However, the example of the Boltzmann brain problem demonstrates that this methodological ignorance cannot be sustained indefinitely even within physics: it conflicts with cosmologists’ goal to say as much as possible about which models of the universe are plausible.
@@
Our work suggests several interesting directions for further research. First, we have not said much about the interpretation of probability theory that would be sufficient or necessary to provide a conceptual foundation of our argumentation. Myrvold’s notion of “epistemic chance” [75] may, for example, constitute an adequate underpinning, and it would be interesting to explore this conjecture in detail. Second, our perspective via Restriction A may be seen as a motivation to explore alternative approaches in the foundations of physics which are, broadly speaking, “idealist” in nature – beginning with observations (and their probabilities) as primary notions, and yielding intersubjective events as approximate, emergent features. One of us has initiated an approach of this kind [63], and adherents to the interpretation of quantum theory called QBism [58] might be sympathetic to attempts of this kind, at least if they can somehow make peace with a non-fully-subjectivist interpretation of probability.
@@
Besides the question of what to make of Restriction A in itself, we hope that our work can inspire some novel connections between puzzles that have so far been studied separately in the philosophy and quantum physics communities. We believe that a unified study of this feature can lead to interesting insights that could not be obtained by considering any one of its instances in isolation.

### v2

@@ III.1 Definition of Restriction 1
Proponents of the Lewisian ‘worm view’, on the other hand, might say that the type of uncertainty that, for example, Freya faces at the beginning of Thought Experiment II.2 is nothing but instantaneous self-locating uncertainty: Freya does not know which spacetime worm she is (the one passing through the green room, or the other one passing through the blue room at future spacetime points). At least in the context of classical physics, Lewisians might argue that their view suggests a natural ‘division of labour’: epistemology or the philosophy of mind would be the fields to consult if Freya would like to obtain guidance in the face of her self-locating uncertainty, whereas physics can inform her (and us) about which types of spacetime worms exist, or which ones are probable, according to the laws of (statistical) mechanics. Restriction 1 would then be an unavoidable consequence for all theories that we deem physical. However, this conclusion becomes problematic in light of quantum theory: according to some Everettian interpretations, quantum probabilities are of a similar kind to the self-locating uncertainty probabilities (according to Lewisians) in branching scenarios like Thought Experiment II.2, and yet, quantum theory as a physical theory has certainly something to say about those probabilities.
@@
In Subsection III.3, we will argue that Restriction 1 can be considered an essential ingredient of another puzzle in the foundations of physics and philosophy: the Boltzmann brain problem. Before we turn to this, let us discuss in more detail how Restriction 1 is related to Bell nonlocality in quantum theory.
III.2 Quantum physics: lifting all Bell inequality violations to instances of Restriction 1
Consider a Bell scenario with associated probabilities $P(a,b|x,y)$, where $x,y$ are Alice’s and Bob’s settings, and $a,b$ are their outcomes. Due to Fine’s Theorem [69] (see also [70]), if some Bell inequality is violated, then there does not exist a joint probability distribution $P(a_{1},\ldots,a_{M_{A}},b_{1},\ldots,b_{M_{B}})$ which reproduces $P(a,b|x,y)$ via marginalization, where $a_{j}$ denotes Alice’s outcome on setting $j$, and $M_{A}$, $M_{B}$ are the numbers of Alice’s and Bob’s possible settings respectively. Hence, the physical properties $a_{j},b_{k}$ do not have a joint probabilistic description. We can understand the main idea of Bong et al.’s EWFS scenarios as lifting the nonexistence of these joint distributions (which are over potentially measured outcomes) to instances of Restrictions A (which is about actually observed outcomes).
@@ III.2 Quantum physics: lifting all Bell inequality violations to instances of Restriction 1
Violating an LF inequality is a mathematically strictly stronger fact (and hence a metaphysically strictly more dramatic fact) than the violation of a Bell inequality. However, it is instructive to consider the results of [71], who have shown that all Bell violations can be lifted to some kind of LF violation. In scenarios where we can identify agents with physical subsystems, there is hence a close relation between Bell nonlocality and Restriction 1 in quantum theory. However, as we have discussed in the context of Definition 2, and and as is evident in Thought Experiments II.2 and II.3, we should not assume that this identification holds in all cases. The result of [71] is shown by considering sequential EWFSs. A sequential EWFS is a scenario in which the superobserving parties (in this case, just Alice) may unitarily reverse the lab and ask their friend to measure in a different basis multiple times prior to ending the experiment. After each measurement performed by Charlie, Alice chooses one of two settings $x_{i}\in\{0,1\}$, where $1\leq i<R$. If $x_{i}=0$ and $i<R$, she reverses the lab and asks Charlie to perform a new measurement. Or if $x_{i}=1$, she asks Charlie his outcome. For the $R$th iteration, after $R-1$ potential reversals, she measures the particle directly for $x_{R}=0$, or asks Charlie’s outcome for $x_{R}=1$. It is shown that for any two-party scenario $\mathcal{S}=(\mathcal{A},\mathcal{B},\mathcal{X},\mathcal{Y})$, the violation of a Bell inequality within a EWFS implies the violation of a LF inequality, i.e.
@@
where $\mathbb{L}\mathbb{D}$ denotes the polytope describing behaviours that satisfy local determinism (Bell correlations) and $\mathbb{S}\mathbb{W}$ those that satisfy the LF assumptions for a sequential EWFS, with $R=|\mathcal{X}|-1$. We will now look at the precise definition of these sets.
@@
A proof of equation (7) is presented in [71]. The first, more established direction is that probabilities in $\mathbb{L}\mathbb{D}(\mathcal{S})$ must also be contained within $\mathbb{S}\mathbb{W}(\mathcal{S}_{R,0})$. This had already been observed in [7], since the LF set of correlations is a strict superset of the LHV set. An intuition for the opposite direction can be obtained by noticing that AOE” implies both AOE’ and Predetermination’. First, AOE” implies AOE’ for $p(\tilde{a}b|\tilde{x}y)=\sum_{\overleftrightarrow{c}}p(\tilde{a}b\overleftrightarrow{c}|\tilde{x}y)$, when $\tilde{a}$ and $\tilde{x}$ are relabelled $a$ and $x$. Second, AOE” implies Predetermination’, as one can define a hidden variable $\zeta$ containing $\overleftrightarrow{c}$ and $j$, where $j$ labels the extreme points of $\mathbb{N}\mathbb{S}((\mathcal{A},\mathcal{B},R+1,\mathcal{Y}))$. Charlie’s outcomes $\overleftrightarrow{c}$ function as deterministic hidden variables for all but one of the settings chosen by Alice, whilst the distribution over $j$ corresponds to the convex mixture of the different deterministic non-signalling behaviours that are realised in the experiment. Intuitively, it is essentially the claim of the absoluteness of Charlie’s outcomes $\overleftrightarrow{c}$ that lets them play the role of a set of hidden variables. For details, please see the proof of [71].
The equivalence expressed by equation (7) demonstrates that all Bell violations can be lifted to an instance of Restriction 1 for quantum theory via an EWF scenario, when one assumes Local Agency. It is also interesting to note that, whilst the quantum thought experiments above all involve instances of violations of Bell-type inequalities, one could also begin with contextuality, and lift this to examples of Restriction 1. This is done, for example, in Refs. [73, 74].
III.3 Restriction 1 beyond quantum foundations: should you believe you are a Boltzmann brain?
We contend that further puzzles in physics and philosophy concerning identity and first-person experience can be reduced to Restriction 1. In this subsection, we discuss one such example for which we argue that this is the case: the Boltzmann brain (BB) problem. For a brief introduction to this problem, see Appendix V.5. Here, we will be even more brief, and discuss only a schematic version ignoring all details that are irrelevant for our discussion. Our exposition will mainly follow Carroll’s work [75].
@@ III.3 Restriction 1 beyond quantum foundations: should you believe you are a Boltzmann brain?
In this paper, we have carefully avoided discussing situations in which agents wonder who or where they are, and we have defined the notion of Restriction 1 in terms of an agent’s observations. To connect the BB discussion to the one in earlier sections, let us therefore reformulate this statement in a way that makes it more operational:
(*) If Freya lives in a BB-dominated universe, then she should expect that she will probably make a BB observation soon (such as seeing the microwave background being replaced by thermal background).
We argue that this reformulation captures more carefully the empirical claim that underlies the standard argument (S) above: after all, we use (what we think are) our observations to conclude that we are not BBs. Moreover, quantum physics motivates some caution to ground physical or metaphysical claims in concrete observations as far as possible. For example, the question of whether a photon in an interferometer is in the left or the right arm cannot be given any immediate meaning unless it is operationalised in terms of an empirical observation (say, by asking whether a detector in the left arm does or does not click). Therefore, statement (*) is a more careful formulation than the one preceding it.
Statement (*) can also be understood as grounding (C), cognitive instability, in a more operational way: if we have randomly fluctuated into existence, then our future observations will probably be uncorrelated with our beliefs and presumed knowledge. In other words: assuming the cosmological model of a BB-dominated universe, (*) implies that all our future observations would be uncorrelated with all we know, including this model and its predictions, which in turn should undermine our trust in the model.
However, statement (*) is far from obvious. It relies on something like the following unspoken assumption:
@@
However, the statement of whether Freya will soon make a BB observation (more succinctly and imprecisely, whether “she” is a BB) is not a statement about facts of the world. As such, we cannot use our physical theories directly to assign a probability $p$ to it. More concretely, physics itself, i.e. our universe’s laws and mechanisms, cannot determine its truth value, or the chances of it becoming true. Even if there are laws of the universe that can be phrased in statistical language in the regime of approximation that is relevant for all copies of Freya, these laws will in general fail to imply directly an assignment of probability $p$. The vast probability space describing all relevant facts of the world, i.e. whatever classical facts will present themselves to you and your fellow physicists according to your joint choice of Heisenberg cut, will contain $N$ random variables (one for each representation or realisation of Freya), $M$ of them for BBs and $N-M$ of them for non-BBs. But if You, the reader, are “Freya”, defined by your locally available information and nothing else, then you cannot identify any random variable on this space that would describe You. Hence, it is impossible to obtain a marginal distribution from a joint distribution of “facts of the world” that would correspond to the correct probability assignment of what You should believe to observe next. This is an instance of Restriction 1.
We can see the claim above (that one should assign probability $M/N$) as a particularly intuitive attempt to respond to Restriction 1 – that is, to follow Elga’s Principle of Indifference [41] in assigning a uniform probability over all realisations. However, the situation described by Elga (an entertaining scenario involving “Dr. Evil” summarised in Appendix V.4) involves only very few (two) copies of an agent, with the property that the future observations of each copy do not dramatically differ from each other in their information-theoretic properties. It is unclear though whether (or, at the very least, not self-evident that) the uniform distribution is what we should assign in all other cases. In particular, the indifference principle must fail if there are infinitely many copies of Freya. Unrealistic as this may seem, we should not exclude this case a priori, since we would like to construct a principled way to respond to Restriction 1 that would also work in a literally infinite universe. Furthermore, Everettians in particular may want to admit that (branch or copy) counting is not always what one should do, but other structural features should impact probability assignments [76]. Before describing an example of what could potentially replace the indifference principle, let us acknowledge that the above statement (*) relies crucially on this assumption, and hence:
@@
We cannot argue that the prediction of a Boltzmann Brain-dominated universe invalidates a cosmological model, unless we respond to Restriction 1 in a particular way [^3] The notion of “agent” is here somewhat different from that of the previous thought experiments, but still in accordance with Definition 2. Namely, in our thought experiments and in WF-type experiments, we can at least in its initial stage identify the agent with a fixed, uniquely identified material object. Here, however, the agent is defined as a user of a physical theory that happens to be in a specific local configuration, without specifying which specific piece of matter corresponds to it, or whether (for example) it is itself a Boltzmann brain or not. Still, in line with Definition 2, every reader can in principle imagine putting themselves in the shoes of this element..
@@
This is the main point we would like to make in this section: Restriction 1 is relevant beyond the thought experiment of Wigner’s Friend. Its reappearance in other scenarios with direct relevance to physics should be regarded as an argument against an attitude of methodological ignorance or of complete deference to other fields of inquiry such as the Philosophy of Mind.
@@
Now, as demonstrated in [65], if the above prescription is assumed to define the chances of Freya’s future observations, then this particular way to respond to Restriction 1 will lead to a consequence that is in some sense the opposite to the one drawn via indifference: it implies that Freya should not expect to make BB-observations soon, even if she lives in a BB-dominated universe. In a nutshell, the reason for this is that the BB-observation of thermal, uncorrelated radiation makes the resulting data algorithmically uncorrelated with Freya’s previous observations, which makes $M(y|x)$ small — crucially, this value becomes much larger if $y$ describes “business as usual on Earth”. We refer the interested reader to [65] for details. This demonstrates the sensitivity of the usual BB argumentation to the choice of “self-location” measure, i.e. to the way in which we respond to Restriction 1.
@@ IV Conclusions
In this article, we have presented a property “Restriction 1”, which we argue is a feature of classical, quantum and indeed more general physical theories: essentially, that a physical theory cannot always provide a probabilistic description of the observations of all agents. We argue that the violation of a Local Friendliness inequality in Extended Wigner’s Friend (EWF) scenarios proves a particularly dramatic instance of Restriction 1 for quantum theory, and all empirically adequate future theories, if one assumes Locality and No Superdeterminism, but that this should be seen amongst a broader class of puzzles that demonstrate the feature. In particular, we have presented a number of thought experiments involving the duplication of agents, for which classical physics provides no means of assigning outcome probabilities for the users of the theory. Moreover, even supplementing classical physics with some probability rule that supplies these predictions will inevitably lead to inconsistent predictions amongst agents (i.e. to a theory that is still subject to Restriction 1), unless this probability rule disregards elements of the setup that are at least intuitively essential. Furthermore, we have shown that Restriction 1 is also at the heart of cosmology’s Boltzmann brain (BB) problem. In particular, the question of whether cosmological models that predict a BB-dominated universe should be disregarded a priori cannot be answered without reacting to Restriction 1 in a particular way.
@@
We take it that the ultimate importance of EWF and similar thought experiments lie in revealing a crucial methodological restriction of our current physical theories: that they typically only provide (probabilistic) predictions for situations in which these predictions can be intersubjectively tested by external observers. However, there are situations such as EWF scenarios for which potential predictions could not be intersubjectively, but privately tested: we could certainly imagine putting ourselves into the shoes of the Friend and testing our predictions for what we will observe, but our theories do not always tell us what to expect in situations of this kind. The philosopher Thomas Nagel has famously asked “What Is It Like to Be a Bat?” [78], emphasizing the significance of the “subjective character of experience”. In the context of EWF scenarios, we can analogously ask: What is it like to be a Friend? In analogy with Nagel’s argument, our claim here is that external facts supplied by our physical theories cannot provide a complete picture of internal experiences and predictions. Crucially, this observation does not only refer to the obvious restriction of not knowing what it feels like to be a Friend, but even of asking whether the Friend will hear a detector click or not. It may be tempting to ignore such questions as “unphysical”, since their answers cannot be verified by experiments performed by external observers. However, the example of the Boltzmann brain problem demonstrates that this methodological ignorance cannot be sustained indefinitely even within physics: it conflicts with cosmologists’ goal to say as much as possible about which models of the universe are plausible. Moreover, to dismiss questions that we may wish to ask, concerning what we ourselves will observe next, as being fundamentally nonsensical seems to deny their genuine relevance for human experience [79].
@@
Our work suggests several interesting directions for further research. First, we have not said much about the interpretation of probability theory that would be sufficient or necessary to provide a conceptual foundation of our argumentation. Myrvold’s notion of “epistemic chance” [80] may, for example, constitute an adequate underpinning, and it would be interesting to explore this conjecture in detail. Second, our analysis via Restriction 1 challenges Elga’s Principle of Indifference (see Theorem 1), and perhaps (in line with Adlam [81]) the received view on self-locating credences altogether. This conclusion could be taken as motivation to explore alternative approaches in the foundations of physics that are, broadly speaking, “idealist” in nature – beginning with observations (and their probabilities) as primary notions, and yielding intersubjective events as approximate, emergent features [65, 79].
@@
Besides the question of what to make of Restriction 1 in itself, we hope that our work can inspire some novel connections between puzzles that have so far been studied separately in the philosophy and quantum physics communities. We believe that a unified study of this feature can lead to interesting insights that could not be obtained by considering any one of its instances in isolation.